1g of 2 oxide of metal are reduced by H2. Metal formed weigh 0.888 and 0.799.show that these data illustrate law of multiple proportion. 
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Let the two oxides of Cu be named A and B 
oxide A
mass of Cu in 1g of oxide A = 0.888g  
mass of O in 1g of oxide A =1-0.888=0.112g
oxide B 
mass of Cu in 1g of   oxide B = 0.799g
mass of O in g of oxide B = 1-0.799=0.201g
In oxide A , 1g Cu is combined with oxygen = 0.112 x 1/0.888 = 0.126g
In oxide B , 1g Cu is combined with oxygen = 0.201 x 1/0.799 = 0.252g
the different weight of oxygen combined with a fixed weight (1g) of Cu are in ratio 0.126 : 0.252  or 1: 2 
This is a simple whole ratio that shows the agree with the law of multiple properties . 
 
 
Regards
Arun (askIITians forum expert)